# Area Selective Chemical Vapor Deposition of Gold by Electron Beam Seeding Zusammenfassung

### v1

@@ Introduction
Electron Beam Seeding
@@
All experiments were carried out in a conventional dual beam instrument fitted with a gas injection system to introduce precursor molecules locally and a sample heating stage, as detailed in the Supporting Information (SI). The autocatalytically enhanced CVD process was found to work reliably on all used bulk substrates. Nevertheless, it is crucial to clean the sample surface prior to the process. Any organic or inorganic residues on the surface hinder the reproducibility of the process by partially impeding autocatalysis, or by functioning as preferred precursor chemisorption sites and thus "contamination seeds."
@@
The precursor compound used, Au(acac)Me; dimethyl(acetylacetonate)gold(III), is primarily known for its use in CVD where it provides pure gold deposition in plasma assisted processing at low temperatures 22. The thermal CVD process was carried out at a substrate temperature of 300 °C. It produced gold films of medium quality with 85 at.% gold content. However, the corresponding fluorinated diketonates, namely tfac and hfac (tetra- and hexafluoroacetonate), resulted in pure gold deposition 23. Au(acac)Me is also commonly utilised in focused electron beam induced deposition (FEBID), a direct writing technique where the precursor molecules physisorbed to the surface are dissociated using a focused electron beam 18.
@@
As electron-induced dissociation is not bonding selective and occurs at room temperature, the use of metal-organic precursor compounds often results in a carbon-rich deposit 24. Efforts have been made to deposit materials with inorganic ligands, which have shown promise especially for magnetic materials 25. For gold, the inorganic precursor, Au(CO)Cl, containing both chlorine and carbonyl ligands, allows high-purity deposition using focused electron beam writing. However, the chlorine content restricts spatial selectivity due to etching effects 26. Similar to PFAuCl, which also provides high-purity gold deposition 27, this compound is not stable under vacuum conditions, severely limiting its practical usability 28. It would be of significant interest to explore gold carbonyls in this context. Although they can be synthesised, gold carbonyls exist only in specific laboratory environments and are therefore not suitable for typical FEBID or CVD experiments 29, 30. For this reason, precursor compounds with organic ligands were used in most of the reported FEBID experiments 31.
@@
The key advantage of the direct electron beam writing process is that it provides extreme shape flexibility in three dimensions 32. To efficiently utilize these distinctive 3D capabilities, several techniques have been employed to attain pure deposits. Here, we will particularly focus on the influence of heat, either as beam-induced heating, especially for large beam currents and thin 3D geometries, or applied externally by sample heating at the various stages of processing. Heat input is viewed as a means of enhancing deposit purity as it promotes the desorption rates of cleaved ligands and contributes to ligand dissociation by thermal assistance 31. However, the acac ligand has a low desorption rate and a tendency to polymerize under electron beam impact. This limits the achievable gold contents to approximately 30 at.% when using thermal energy input 33, 34. To achieve pure 3D gold structures, additional oxidizing agents must be used in combination with elaborate shape optimization for effective purification 35, 36, 37. The key question is whether this presents a fundamental constraint, or whether thermal assistance at low electron fluxes, as in the case of silver deposition 38, 39, can result in pure yet still selective metal deposition. In this scenario, the clean surface of the gold seed could even be autocatalytically active.
@@
To investigate this hypothesis, the following experimental matrix was performed. Under a continuous gas flow, an array of 2x2 µm squares was patterned with varying electron beam doses using a beam energy of 5 keV and a beam current of 100 pA. Figure 2(a) shows part of the array, the complete array is available in the SI. As a representation of the applied electron dose, the patterning time is provided in milliseconds within each square. The corresponding pixel spacing (pitch) and number of repetitions (repeats) are also indicated alongside. The utilized pitches of 10 and 5 nm correspond to doses of 6.25 and 25 electrons per nm and repeat, respectively. Starting from a single repeat which is supposed to deposit only isolated gold atoms and small clusters up to 100 repeats were conducted, as the latter would lead to a thin FEBID deposit. To investigate the effect of ligand desorption, the seed deposition was carried out at room temperature (30 °C) and at elevated stage temperatures. Similar to early reports on Fe(CO) by Kunz 12, we observed efficient electron-enhanced CVD through continued electron irradiation under low fluxes and elevated stage temperatures (see Figure SI.3). Therefore special care was taken to avoid unintended beam exposure, and no imaging was conducted before or after the seed array deposition. As the GIS system has an automatically controlled valve, gas flow occurred only during seeding for the array deposition time of 1 min 18 s. As a control experiment for electron beam induced surface activation, the same array was deposited at the different substrate temperatures without gas flow (see Figure SI.4).
Autocatalytic Growth of Gold
@@
In order to achieve autocatalytic activity of the seed layer, two conditions must be met. Firstly, the previously deposited metallic nuclei must have a clean surface to participate in the reaction. Therefore, ultra-high vacuum (UHV) conditions are commonly used for this type of experiment. Additionally, all final reaction products that result from ligand cleavage must be volatile enough to desorb intactly 31. Secondly, the growth kinetics at the nucleus must promote chemisorption of the precursor molecules through the removal of the initial ligand with activation energy lower than the standard bond dissociation energies 15.
This condition applies to some carbonyls, e.g. area selective CVD by electron beam seeding has been demonstrated for the deposition of transition metals using precursors like Fe(CO) 12, 21, Co(CO)NO 20, and Cr(CO) 15. In some cases high vacuum (HV) conditions were sufficient to observe autocatalysis, e.g. for Fe(CO) 40, but in others clean metal surfaces were additionally required excluding electron beam writing for seeding, e.g. for Co(CO) 41. Under ultra high vacuum conditions, NH dosing was found to inhibit autocatalytic growth, which could subsequently be reactivated by electron beam seeding, providing an additional control option in area selective CVD utilising Fe(CO) 21. Generally, carbonyls possess the thermodynamically stable CO ligand, making them the preferable choice for autocatalytically enhanced growth processes. In contrast, organic ligands such as the frequently used -diketonates (here acac) are not expected to have such favorable desorption properties as mentioned above.
@@
The autocatalytic growth experiments were performed in the same vacuum chamber. After depositing a square array for one temperature, the electron beam was switched off and the GIS nozzle was opened again for 30 min to deliver precursor molecules.
@@
The growth results are shown in Figure 2(b) where carbon appears as a dark contrast and gold as a bright contrast, which is a typical characteristic of SE imaging. At 60 °C, no observable difference was noted when compared to the seed arrays deposited initially. However, at 70 °C, initial signs of CVD growth became evident. The contrast shifted from dark (carbon-rich) to bright (gold-rich), becoming more pronounced as the growth temperature increased. Squares with a greater number of repeats in the electron beam seeding are brighter, corresponding to more pronounced gold growth. This can be attributed to an increased density of gold nuclei available with deposition time. The optimal balance between growth rate and selectivity is achieved at a temperature of 80 °C. Raising the temperature further enhances the growth rate, but decreases selectivity due to the initiation of growth in the halo regions around the actual patterned squares. This phenomenon is demonstrated in Figure 2(c), a high-resolution SE micrograph of the square seeded at 90 °C with 20 repetitions at 5 nm pitch. Although there are no primary electrons in the halo, there are still secondary and backscattered electrons present with a significantly lower flux than in the deposition region 38. Consequently, numerous repetitions in seeding at 80 °C can also diminish the precision of square edges.
@@
Interestingly, electron beam seeding also works at room temperature. However, as anticipated, the higher desorption rate of ligands during seeding at elevated stage temperature suppresses carbon co-deposition, resulting in higher densities of catalytically active seeds. This is illustrated in Figures 1(a) and (b). Therefore, the films grown from seed layers deposited at room temperature show a significant carbon contrast, while at the elevated stage temperature the seed density is already high enough to form a quasi-continuous gold film, cf. Figure 2(b) middle row.
@@
For the tested pitch values and numbers of repeats, the growth of the autocatalytic film at a certain temperature is only influenced by the total electron dose. When comparing SE contrast and topography, it is evident that the seeds, which were deposited with different pitches and repeats but the same total dose, display the same thickness and SE contrast (grain size) after growth. The optimum dose is reached at about 140 pC/µm or 875 electrons per nm. Additionally, the process exhibits exceptional selectivity as gold growth is restricted to the presence of gold nuclei under these temperatures. It is worth noting, however, that a substantial decrease in CVD growth rate was observed over time, resulting in saturation for attainable thicknesses at approximately 6.5 nm. This could be due to surface poisoning, i.e. carbon coverage of the active gold nuclei due to residual hydrocarbons in the vacuum chamber and/or low desorption rates of the ligand fragments. In this instance, an improvement in the background vacuum would be promising. Another potential reason is the nuclei changing in size during autocatalytic growth. Size effects can have a significant impact on the catalytic activity of metallic nanoparticles, especially in the case of gold 42. It is worth mentioning that the observed decreasing CVD growth rate is not an obstacle to technological implementation, as seeding and growth can be easily iterated several times (as shown in Figure SI.3).
@@
Beam-induced Heating and Autocatalysis
@@
One of the main benefits of using a tightly focused electron beam to decompose molecules from the gas phase is its extreme 3D flexibility, resolution and shape fidelity 32. Generally, FEBID is considered to be stimulated by secondary electrons which dissociate the physisorbed precursor molecules. Electron-induced chemistry involves various reaction pathways, which depend on the local availability of electrons, molecules, and co-reactants 25, 31. Beam-induced heating is also a significant factor. As the growth surface locally heats up, the dynamics of the supplied precursor molecules and fragments will shift, e.g. towards faster diffusion and higher desorption rates. Estimates show that even slight changes in surface temperature, around 10 K, can significantly impact the vertical growth rate during 3D FEBID 43, 44, 45, 46, which typically operates in the reaction-rate limited growth regime. However, there is no experimental data available regarding the actual local temperature changes that may occur. Here, we exploit the onset of autocatalytic CVD during the growth of one-dimensional vertical nanostructures (pillars) under steady-state growth conditions as a means of in situ temperature measurement.
@@
Figure 4(a)-(c) shows example images of deposited nanopillars using the precursor Au(acac)Me with diameters of about 50 nm and aspect ratios of 40:1 and larger. The primary energy used was 15 keV at a beam current of 100 pA, which aligns with typical standards in 3D FEBID printing. Stationary growth conditions were achieved using a dwell time of 26.2 ms 43, 47. Each pillar’s overall growth duration was consistently maintained at 2 min 30 s deposition time.
@@
Figure 4 (a) presents the nanopillar obtained at room temperature, showing clear bright particles in the upper part. The presence of gold particles confirms the existence of autocatalytically active gold nuclei on the surface of the pillar and the attainment of the temperature threshold T for visible particle growth. By adjusting the substrate temperature, the region of particle growth onset can be moved along the pillar, as demonstrated by the white arrows in Figures 4(a)-(c). The corresponding substrate temperatures are subsequently plotted as a function of the vertical distance from the substrate surface to the onset region, cf. Figure 4(d). The resulting decrease in with the increase in substrate temperature is linear.
@@
This linear relationship is in line with the fact that there is a negligible loss of thermal energy across the surface of the pillar (refer to SI for estimates on radiative and convective heat losses). Therefore, all the thermal energy introduced by the stationary electron beam is conducted through the pillar into the substrate, which can be considered as a constant temperature heat sink. The heating induced by the beam during growth results in a linear temperature gradient within the pillar 18. Therefore, the vertical position of the CVD onset region is dependent on the stage temperature in a linear manner. The overall height of the pillars reduces for higher stage temperatures, as a result of the increased desorption rate of precursor molecules.
@@
The intersection point of the linear fit with the temperature axis yields a measured temperature of T 112 °C, which indicates the point at which particle growth becomes visible. The thermal decomposition of the precursor conforms to Arrhenius’ law, with the autocatalysis rate exponentially dependent on temperature , where the activation energy acts as the threshold value and is the Boltzmann constant. The overall growth duration was relatively short, requiring a significant growth rate to observe particle growth. These findings align with our observations of the electron beam-seeded autocatalytic growth at temperatures of 100-110 °C, which was highly efficient but showed decreasing selectivity. For 3D growth, reduced selectivity is not an issue since only the thermally decoupled areas reach high enough temperatures.
@@
The proportionality constant K/nm indicates the temperature variation per unit length along the pillar. It offers a quantitative measure of the thermal energy introduced by the electron beam, providing valuable insights for modelling 3D electron beam printing and the possibility of quantitatively determining the thermal conductivity of the deposited material. For this purpose, the measurements are supported by Monte-Carlo modelling of the electron trajectories obtained for this geometry. A mass density range of gcm is assumed, corresponding to the typically observed Au contents in deposits of this precursor, cf. SI for more details on the Monte-Carlo modelling. The obtained range for the average inelastic energy loss keV per incident electron inside the nanostructure can then be used to determine the thermal conductivity of the deposit material:
@@
with the electron beam current , elementary charge and the cross-sectional area of the nanopillar. This can be considered an upper limit since some inelastic losses are due to the production of secondary electrons (and X-rays), while the rest results in Joule heating (e.g. from an assessment of this geometry 44).
@@
The value 0.4 W/(Km) obtained for thermal conductivity has an uncertainty margin of approximately 24 % and falls within the commonly observed range for various carbon modifications, particularly those ranging from amorphous to diamond-like carbon 48. By repeating this type of experiment for different acceleration voltages, beam currents and possibly other precursors, the thermal conductivity of the FEBID material for the different deposition conditions can be determined quantitatively.
Conclusions

### v2

@@ 2 Results and Discussion
2.1 Electron Beam Seeding
@@
All experiments were carried out in a conventional dual beam instrument fitted with a gas injection system (GIS) to introduce precursor molecules locally and a sample heating stage, as detailed in the Supporting Information (SI). The autocatalytically enhanced CVD process was found to work reliably on all used bulk substrates. Nevertheless, it is crucial to clean the sample surface prior to the process. Any organic or inorganic residues on the surface hinder the reproducibility of the process by partially impeding autocatalysis, or by functioning as preferred precursor chemisorption sites and thus ”contamination seeds.”
@@
The precursor compound used, Au(acac)Me2; dimethyl(acetylacetonate)gold(III), is primarily known for its use in CVD where it provides pure gold deposition in plasma assisted processing at low temperatures [22]. The thermal CVD process was carried out at a substrate temperature of 300 °C. It produced gold films of medium quality with 85 at.% gold content. However, the corresponding fluorinated diketonates, namely tfac and hfac (tetra- and hexafluoroacetonate), resulted in pure gold deposition [23]. Au(acac)Me2 is also commonly utilized in focused electron beam induced deposition (FEBID), a direct writing technique where the precursor molecules physisorbed to the surface are dissociated using a focused electron beam [18].
@@
Electron-induced dissociation is mainly triggered by secondary electrons with an energy distribution between 0 and 50 eV. Therefore, direct electron beam writing is not bond selective and the use of metal-organic precursor compounds often results in a carbon-rich deposit [18, 24]. Efforts have been made to deposit materials with inorganic ligands, which have shown promise especially for magnetic materials [25]. For gold, the inorganic precursor, Au(CO)Cl, containing both chlorine and carbonyl ligands, allows high-purity deposition using focused electron beam writing. However, the chlorine content restricts spatial selectivity due to etching effects [26]. Similar to PF3AuCl, which also provides high-purity gold deposition [27], Au(CO)Cl is not stable under vacuum conditions, severely limiting its practical usability [28]. It would be of significant interest to explore gold carbonyls in this context. Although they can be synthesized, gold carbonyls exist only in specific laboratory environments and are therefore not suitable for typical FEBID or CVD experiments [29, 30]. For this reason, precursor compounds with organic ligands were used in most of the reported FEBID experiments [31].
@@
The key advantage of the direct electron beam writing process is that it provides extreme shape flexibility in three dimensions [32]. To efficiently utilize these distinctive 3D capabilities, several techniques have been employed to attain pure deposits. Here, we will particularly focus on the influence of heat, either as beam-induced heating, especially for large beam currents and thin 3D geometries, or applied externally by sample heating at the various stages of processing. Heat input is viewed as a means of enhancing deposit purity as it promotes the desorption rates of cleaved ligands and contributes to ligand dissociation by thermal assistance [31]. However, the acac ligand has a low desorption rate and a tendency to polymerize under electron beam impact. This limits the achievable gold contents to approximately 30 at.% when using thermal energy input [33, 34]. To achieve pure 3D gold structures, additional oxidizing agents must be used in combination with elaborate shape optimization for effective purification [35, 36, 37]. The key question is whether this presents a fundamental constraint, or whether thermal assistance at low electron fluxes, as in the case of silver deposition [38, 39], can result in pure yet still selective metal deposition. In this scenario, the clean surface of the gold seed could even be autocatalytically active.
@@
To investigate this hypothesis, the following experimental matrix was performed. Under a continuous gas flow, an array of 2x2 µm2 squares was patterned with varying electron beam doses using a beam energy of 5 keV and a beam current of 100 pA. Figure 2(a) shows part of the array, the complete array is available in the SI. The applied electron dose in pC/µm2 is provided within each square. The corresponding pixel spacing (pitch) and number of repetitions (repeats) are also indicated alongside. The utilized pitches of 10 and 5 nm correspond to doses of 6.25 and 25 electrons per nm2 and frame repeat, respectively. Starting from a single repeat which is supposed to deposit only isolated gold atoms and small clusters up to 100 repeats were conducted, as the latter would lead to a thin FEBID deposit. In addition, the effective local dose changes when the pitch is varied, as the dwell time is kept constant. With the same total dose, seed layers may exhibit different purities as co-dissociation of ligands or a too slow ligand desorption rate could play a role [40]. To investigate the effect of ligand desorption, the seed deposition was carried out at room temperature (30 °C) and at elevated stage temperatures. Similar to early reports on Fe(CO)5 by Kunz [12], we observed electron-enhanced CVD i.e. an accelerated growth rate through continued electron irradiation under low electron fluxes and elevated stage temperatures (see Figure SI.4). Therefore special care was taken to avoid unintended beam exposure, and no imaging was conducted before or after the seed array deposition. As the GIS system has an automatically controlled valve, gas flow occurred only during seeding for the array deposition time of 1 min 18 s. Different control experiments for electron beam induced surface activation and for seeding using other precursors were performed with the same array deposited at the different substrate temperatures (see Figures SI.5-.7).
2.2 Autocatalytic Growth of Gold
In order to achieve autocatalytic activity of the seed layer, two conditions must be met. Firstly, the previously deposited metallic nuclei must have a clean surface to participate in the reaction.
@@ 2.2 Autocatalytic Growth of Gold
Additionally, all final reaction products that result from ligand cleavage must be volatile enough to desorb intactly [31]. Secondly, the growth kinetics at the nucleus must promote chemisorption of the precursor molecules through the removal of its ligands [15].
This condition applies to some metal carbonyls, e.g. area selective CVD by electron beam seeding has been demonstrated for the deposition of transition metals using precursors like Fe(CO)5 [12, 21],
Co(CO)3NO [20], and Cr(CO)6 [15].
In some cases high vacuum (HV) conditions were sufficient to observe autocatalysis, e.g. for Fe(CO)5 [41], but in others clean metal surfaces were additionally required excluding electron beam writing for seeding, e.g. for Co2(CO)8 [42].
Under ultra high vacuum conditions, NH3 dosing was found to inhibit autocatalytic growth, which could subsequently be reactivated by electron beam seeding, providing an additional control option in area selective CVD utilising Fe(CO)5 [21].
As the carbonyl ligand is thermodynamically stable and neutral, carbonyls are the preferable choice for autocatalytically enhanced growth processes.
However, as discussed above gold carbonyl precursors are not vacuum stable.
Gold precursors typically contain non-neutral organic ligands such as the frequently used -diketonates (here acac) which have less favorable desorption properties [43].
@@
The autocatalytic growth experiments were performed in the same vacuum chamber. After depositing a square array for one given temperature, the electron beam was switched off and the GIS nozzle was opened again for 30 min to deliver precursor molecules.
@@
The growth results are shown in Figure 2(b) where carbon appears as a dark contrast and gold as a bright contrast, which is a typical characteristic of SE imaging. At 60 °C, no observable difference was noted when compared to the seed arrays deposited initially. However, at 70 °C, initial signs of CVD growth became evident. The contrast shifted from dark (carbon-rich) to bright (gold-rich), becoming more pronounced as the growth temperature increased. Squares with a greater number of repeats in the electron beam seeding are brighter, corresponding to more pronounced gold growth. This can be attributed to an increased density of gold nuclei available with deposition time. The optimal balance between growth rate and selectivity is achieved at a temperature of 80 °C, as shown in a close-up Fig. 2(c). Raising the temperature further enhances the growth rate, but decreases selectivity due to the initiation of growth in the halo regions around the actual patterned squares. This phenomenon is demonstrated in Figure 2(d), a high-resolution SE micrograph of the square seeded at 90 °C with 20 repetitions at 5 nm pitch. Although there are no primary electrons in the halo, there are still secondary and backscattered electrons present with a significantly lower flux than in the deposition region [38]. Consequently, numerous repetitions in seeding at 80 °C can also diminish the precision of square edges.
@@
Interestingly, electron beam seeding also works at room temperature. However, as anticipated from [34], the higher desorption rate of ligands during seeding at elevated stage temperature suppresses carbon co-deposition, resulting in higher densities of catalytically active seeds. This is illustrated in Figures 1(a) and (b). Therefore, the films grown from seed layers deposited at room temperature show carbon as low contrast areas, while at the elevated stage temperature the seed density is already high enough to form a quasi-continuous gold film, cf. Figure 2(b) middle row.
For the tested pitch values and numbers of repeats, the growth of the autocatalytic film at a certain temperature is only influenced by the total electron dose. When comparing SE contrast and topography, it is evident that the seeds, which were deposited with different pitches and repeats but the same total dose, display the same thickness and SE contrast (grain size) after growth, cf. Fig. 2(b) for 20 pC/µm2 and SI for more examples. The optimum dose is reached at about 140 pC/µm2 or 875 electrons per nm2.
@@
Additionally, the process exhibits exceptional selectivity as gold growth is restricted to the presence of gold nuclei under these temperatures. It is worth noting, however, that a substantial decrease in CVD growth rate was observed over time, resulting in saturation for attainable thicknesses at approximately 6.5 nm (cf. SI for more details). This could be due to surface poisoning [42], i.e. carbon coverage of the active gold nuclei due to residual hydrocarbons in the vacuum chamber and/or low desorption rates of the ligand fragments. In this instance, an improvement in the background vacuum would be promising. Another potential reason is the nuclei changing in size during autocatalytic growth. Size effects can have a significant impact on the catalytic activity of metallic nanoparticles, especially in the case of gold [44]. It is worth mentioning that the observed decreasing CVD growth rate is not an obstacle to technological implementation, as seeding and growth can be easily iterated several times (as shown in Figure SI.4).
@@
2.3 Beam-induced Heating and Autocatalysis
@@
One of the main benefits of using a tightly focused electron beam to decompose molecules from the gas phase is its extreme 3D flexibility, resolution and shape fidelity [32]. Generally, FEBID is considered to be stimulated by secondary electrons which dissociate the physisorbed precursor molecules. Electron-induced chemistry involves various reaction pathways, which depend on the local availability of electrons, molecules, and co-reactants [25, 31]. Beam-induced heating is also a significant factor. As the growth surface locally heats up, the dynamics of the supplied precursor molecules and fragments will shift, e.g. towards faster diffusion and higher desorption rates. Estimates show that even slight changes in surface temperature, around 10 K, can significantly impact the vertical growth rate during 3D FEBID [45, 46, 47, 48]. However, there is no experimental data available regarding the actual local temperature changes at the growing nanopillar apexes where the e-beam impacts. Here, we exploit the onset of autocatalytic CVD during the growth of one-dimensional vertical nanostructures (pillars) under steady-state growth conditions as a means of in situ temperature measurement.
@@
Figure 4(a)-(c) shows example images of deposited nanopillars using the precursor Au(acac)Me2 with diameters of about 50 nm and aspect ratios of 40:1 and larger. The primary electron energy used was 15 keV at a beam current of 100 pA, which aligns with typical standards in 3D FEBID printing [47, 49, 50]. Stationary growth conditions were achieved using a dwell time of 26.2 ms [45, 49]. Each pillar’s overall growth duration was consistently maintained at 2 min 30 s deposition time.
@@
Figure 4 (a) presents the nanopillar obtained at room temperature. It shows a base cylinder with smooth surface and bright contrast gold particles in the upper part. The presence of gold particles confirms the existence of autocatalytically active gold nuclei on the surface of the pillar and the attainment of the temperature threshold Ti for visible particle growth. No CVD growth was observed on the remaining nanopillar surface below and the surface retained its original smooth appearance. By adjusting the substrate temperature, the region of particle growth onset can be moved along the pillar, as demonstrated by the white arrows in Figures 4(a)-(c). The corresponding substrate temperatures are subsequently plotted as a function of the vertical distance from the substrate surface to the onset region, cf. Figure 4(d). The resulting decrease in with the increase in substrate temperature is linear.
@@
This linear relationship is in line with the fact that there is a negligible loss of thermal energy across the surface of the pillar (cf. SI for estimates on radiative and convective heat losses). Therefore, all the thermal energy introduced by the stationary electron beam is conducted through the pillar into the substrate, which can be considered as a constant temperature heat sink. The heating induced by the beam during growth results in a linear temperature gradient within the pillar [18]. Therefore, the vertical position of the CVD onset region is dependent on the stage temperature in a linear manner. The overall height of the pillars reduces for higher stage temperatures, as a result of the increased desorption rate of precursor molecules.
@@
The intersection point of the linear fit with the temperature axis yields a measured temperature of T 112 °C, which indicates the point at which particle growth becomes visible. The thermal decomposition of the precursor conforms to Arrhenius’ law, with the autocatalysis rate exponentially dependent on temperature , where the activation energy acts as the threshold value and is the Boltzmann constant. The overall growth duration was relatively short, requiring a significant growth rate to observe particle growth. These findings align with our observations of the electron beam-seeded autocatalytic growth at temperatures of 100-110 °C, which was highly efficient but showed decreasing selectivity. For growth of high aspect-ratio structures, reduced selectivity is not an issue since only the thermally decoupled areas reach high enough temperatures.
@@
The proportionality constant K/nm indicates the temperature variation per unit length along the pillar. It offers a quantitative measure of the thermal energy introduced by the electron beam, providing valuable insights for modelling 3D electron beam printing and the possibility of quantitatively determining the thermal conductivity of the deposited material. For this purpose, the measurements are supported by Monte-Carlo modelling of the electron trajectories obtained for this geometry. A mass density range of gcm3 is assumed, corresponding to the typically observed Au contents in deposits from this precursor, cf. SI for more details on the Monte-Carlo modelling. The obtained range for the average inelastic energy loss keV per incident electron inside the nanostructure can then be used to determine the thermal conductivity of the deposit material:
@@
with the electron beam current , elementary charge and the cross-sectional area of the nanopillar. This can be considered an upper limit since some inelastic losses are due to the production of secondary electrons (and X-rays), while the rest results in Joule heating (e.g. from an assessment of this geometry [46]).
@@
The value 0.4 W/(Km) obtained for thermal conductivity has an uncertainty margin of approximately 24 % and falls within the commonly observed range for various carbon modifications, particularly those ranging from amorphous to diamond-like carbon [51]. By repeating this type of experiment for different acceleration voltages, beam currents and possibly other precursors, the thermal conductivity of the FEBID material for the different deposition conditions can be determined quantitatively.
3 Conclusions and Outlook